Consider a die with eight sides, marked one, two, three, and so on. Assuming equally likely outcomes, find the probability that the sum of two dice is the given number. (Enter the answer either as a fraction or as a decimal rounded to three places.)
Answer by chessace(471)

Each probability is the count of * ways / 64 possibilities.
Total : Count (you divide by 64)
1 : 0
2 : 1
3 : 2
4 : 3
5 : 4
6 : 5
7 : 6
8 : 7
9 : 8
10: 7
11: 6
12: guess.
...
16: 1
Total : 8 + 2 (sum 1 to 7) = 8 + 2 (28) = 56 + 8 = 64.